Accountancy

Fixed amount is withdrawn on the first day of the month 6.5 . Rajan a partner in a firm withdraws rs 3000 for month. Regularly only the first day of every month. Interest is 12% . Calculate interest on drawings 
Answer #1

Interest on drawings= 3000×12=36000

=36000×6.5/12×100

=2340

answered by: Advika